Handley uses a standard costing system. For production of a fuel cell, standards call for 4 model RN-3 valves per cell costing $2,450 each. During 2018, the company purchased 250 valves for $570,000. The company used 240 valves in the production of 58 fuel cells (8 valves were ruined through installation error).

Calculate material price and quantity variances for the model RN-3 valves. (Enter all variances as a positive number.)

Material Price Variance $

Entry field with incorrect answer

Entry field with incorrect answer Neither Unfavorable nor FavorableUnfavorableFavorable

Material Quantity Variance $

Entry field with incorrect answer

Entry field with incorrect answer UnfavorableNeither Unfavorable nor FavorableFavorable
Answer #1

Materials Price Variance = (SP-AP) *AQ

= (2450 - 570,000/250) * 250

= 42,500 Favourable

.

Materials Quantity Variance = (SQ-AQ) *SP

= (58*4 - 240)*2450

= 19,600 Unfavorable

Comment if you face any issues